# Lin Li1,2∗,  Wei Chen1∗,  Jiahui Li3,  Long Chen1†

### v1

@@ Abstract
Recent advances in multi-modal large language models (MLLMs) have significantly improved object-level grounding and region captioning, but remain limited in visual relation understanding (e.g., scene graph generation), particularly in modeling N-ary relationships that identify multiple semantic roles among an action event. Such a lack of semantic dependencies modeling among multi-entities leads to unreliable outputs, intensifying MLLMs’ hallucinations and over-reliance on language priors. To this end, we propose Relation-R1, the first unified relational comprehension framework that explicitly integrates cognitive chain-of-thought (CoT)-guided Supervised Fine-Tuning (SFT) and Group Relative Policy Optimization (GRPO) within a reinforcement learning (RL) paradigm. Specifically, we first establish foundational reasoning capabilities via SFT, enforcing structured outputs with thinking processes. Then, GRPO is utilized to refine these outputs via multi-reward optimization, prioritizing visual-semantic grounding over language-induced biases, thereby improving generalization capability. Extensive experiments on widely-used PSG and SWiG datasets demonstrate that Relation-R1 achieves state-of-the-art performance in both binary and N-ary relation understanding.
11footnotetext: * The first two authors contribute equally to this work.22footnotetext: † Corresponding Author: Long Chen.
@@ 1 Introduction
Recent advancements in MLLMs have significantly enhanced holistic image understanding and object-level grounding capabilities, enabling tasks like region captioning and referring question answering. However, current MLLMs exhibit critical limitations in relational scene understanding — a core competency for achieving human-like visual cognition, as illustrated in Figure 1(a).
@@
Despite existing studies have made progress in binary relation detection [1, 2], i.e., identifying subject-object interactions (e.g., “child-drinking-glass”), they fail to capture $N$-nary relationship, that identify $N$-th entities engaged in the relation with their roles (what, who, and where concepts). It leads to neglecting the semantic dependencies among multi-entities, such as the glass’s role as the functional container of milk in the drinking action, which may result in unreasonable relation triplets, e.g., “child-drinking-glass”. Furthermore, these unreasonable pair-wise relations may make the MLLMs suffer from both hallucinations and over-relying on language priors rather than visual-semantic evidence. Consequently, there arises an urgent demand for a unified framework capable of addressing these shortcomings. This raises a critical question: How can we devise a unified framework for joint binary and N-ary relational reasoning with grounded cues?
@@
Intuitively, such complicated relational comprehension requires both reasoning and generalization abilities. As for reasoning, the model must perform multi-step and multi-modal reasoning, it includes entity identification, relation recognition, role inference (e.g., “glass” as a container), localization, and contextual integration. In terms of generalization, models trained on limited data struggle to extrapolate to novel entity combinations or ambiguous contexts (e.g., “cup” as a container). In addition, language priors (e.g., common-sense assumptions) often dominate over visual grounding, further limiting their robustness.
Fortunately, DeepSeek-R1-Zero has successfully demonstrated the emergence of reasoning capabilities in LLMs purely through RL. Inspired by this breakthrough, we propose a unified relation comprehension framework Relation-R1, that integrates Supervised Fine-Tuning (SFT) and RL to empower MLLMs with relational reasoning and generalization capability. While SFT alone suffers from poor generalization due to reliance on fixed training patterns, RL struggles to produce structured outputs with consistent formatting [3]. To alleviate these limitations, we design a two-stage pipeline:
Stage 1. SFT: Establishes foundational reasoning through cognitive chain-of-thought guidance. The model generates structured outputs by analyzing visual evidence (e.g., object detection, spatial localization, etc.), ensuring visual grounding over language priors.
Stage 2. RL: Refines outputs via GRPO guided by multi-rewards: (1) format rewards, (2) binary relation rewards, and (3) N-ary relation rewards. This stage employs policy gradient updates to prioritize visual-semantic coherence over language priors, fostering robust relational reasoning and generalization.
By first leveraging SFT to establish reasoning foundations and then applying GRPO to optimize for structured, grounded outputs, Relation-R1 achieves robust relational reasoning and generalization to novel contexts, directly addressing the limitations of prior MLLMs.
@@
We validate the effectiveness of Relation-R1 on both the binary relation detection task (e.g., Scene Graph Generation) and N-ary relation detection task (e.g., Grounded Situation Recognition) under both few-shot and full-data settings. Experiments on PSG [4] dataset and SWiG [5] dataset demonstrate state-of-the-art performance in relational reasoning while maintaining generalization in ambiguous contexts. In summary, our contributions are threefold:
@@
  • We incorporate both binary relation detection and N-ary relation detection in the unified relation comprehension framework.
@@
  • We propose Relation-R1, a unified framework that integrates cognitive CoT guided SFT with RL. To the best of our knowledge, this is the first R1-like work to enable both relational reasoning and robust generalization.
@@
  • Experiments on the prevalent relation understanding datasets under different settings demonstrate the effectiveness of Relation-R1.
@@ 2 Related Work
Reasoning in Multimodal Large Language Models. Recent Large Language Models (LLMs) have demonstrated advanced reasoning capabilities by emulating human-like stepwise thinking, significantly enhancing performance on complex tasks [6, 7]. A pivotal breakthrough, DeepSeek-R1 [7], leverages large-scale reinforcement learning (RL) with formatting and result-oriented rewards to enable LLMs to autonomously generate human-like complex Chain-of-Thought (CoT) reasoning, achieving state-of-the-art results across diverse domains. Multimodal Large Language Models (MLLMs) extend these capabilities through systematic integration of cross-modal knowledge representation and task-specific rewards [8, 9, 10] For multimodal mathematical reasoning, frameworks such as R1-Onevision [8], Virgo [9], and MM-Eureka [11] jointly reason over visual and textual inputs to solve quantitative tasks, while Vision-R1 [10] combines cold-start initialization with GRPO to refine complex CoT reasoning. For fine-grained classification and grounding tasks, Visual-RFT [12] and R1-Omni [13] utilize GRPO-based reinforcement algorithms and verifiable rewards to enhance contextual reasoning. As for the pixel-level understanding task, SegZero [14] employs reinforcement learning to achieve high-resolution semantic segmentation. Despite these advances, existing MLLMs predominantly focus on object-centric recognition rather than unified visual relation understanding, which requires compositional reasoning about geometric, semantic, and functional interactions between objects and contexts. To bridge this gap, our Relation-R1 explicitly integrates relational reasoning into MLLMs via reinforcement learning, enabling systematic inference for both scene interpretation and event prediction.
@@
Grounded Situation Recognition (GSR). GSR extends scene understanding by jointly modeling actions (verbs) and their associated semantic roles. Pratt et al. [5] pioneered this field with a two-stage RNN-based framework: verb detection followed by noun localization. Subsequent works [26, 27, 28, 29] improved this pipeline by integrating transformers and semantic relation modeling, enabling more coherent scene interpretations. Despite these efforts, previous efforts that depend on annotated training samples may face challenges like long-tailed distribution [15, 16, 17], potentially limiting real-world applicability. While LEX [30] realizes training-free zero-shot GSR, it fails to provide an end-to-end framework and exhibits suboptimal performance and computational efficiency due to its reliance on LLMs’ generated descriptions. This paper proposes Relation-R, a unified framework that bridges the gap between open-ended capabilities and end-to-end training.
@@ 3 Methodology
Figure 2: Overview of the proposed Relation-R1 framework. (a) SFT with cognitive CoT, which enforces format alignment through task-specific reasoning processes. (b) RL with GRPO, employs rule-based rewards (e.g., format rewards, binary relation rewards, and N-nary relation rewards) to cultivate robust relation reasoning and generalization ability through policy gradient updates.
@@ 3.1 Preliminaries
Binary Relation Detection. This task requires the model to generate an open-ended scene graph given an image, explicitly marking semantic relationships between pairwise objects [1]. The output is structured as a sequence of relation triplets annotated with formal tags: objects are denoted by <ref> tags followed by their bounding box coordinates < box> to localize the entity in the image, e.g., <ref>person</red></box>[[$x_{1}$, $y_{1}$, $y_{x}$, $y_{2}$]]</box>. Predicates are enclosed in <pred> tags, followed by references to their subject and object entities via their corresponding bounding boxes.
N-ary Relation Detection. Similarly, this task aims to output a open-ended grounded situation frame [5] structured as a primary activity (relation), followed by a sequence of <role> tags along with corresponding bounding box coordinates, each specifying an entity’s role, e.g., <agent>person</agent></box>[$x_{1}$, $y_{1}$, $y_{x}$, $y_{2}$]</box>. Each role explicitly defines the entity’s participation in the activity.
@@ 3.2 Supervised Fine-Tuning
As mentioned above, directly adopting GRPO faces challenges in generating format-inconsistent scene graph captions and grounded situation frame outputs. To this end, we adopt supervised fine-tuning that serves as a foundational stage for task-specific format alignment (c.f., Figure 2(a)).
However, utilizing merely the answer format as the SFT target risks compromising the model’s reasoning capability by encouraging rigid adherence to syntactic structures [3]. To mitigate this, we devise the cognitive Chain-of-Thoughts (CoT), which incorporates stepwise cognitive processes such as object classification, object grounding, visual relationship inference, etc. This ensures the model retains its ability to perform multi-step reasoning while adhering to output formats.
To avoid the MLLM overfitting to specific CoT patterns during SFT, we devise a cognitive CoT generation prompt that explicitly guides an extra strong MLLM (e.g., Qwen 2.5-VL [31]) to generate step-by-step reasoning processes. To be specific, this cognitive CoT generation prompt consists of task definition, scene graph caption/grounded situation frame examples, and GT scene graph caption/grounded situation frame of the current image. The former two make the MLLM generate the correct task format. Meanwhile, the latter one anchors reasoning content to visual evidence, ensuring that the MLLM’s cognitive CoT aligns with both semantic validity and visual grounding rather than relying on category priors. The detailed Cognitive CoT generation prompt is left in the Appendix.
These cognitive CoT are explicitly encapsulated within <think> tags, during SFT. By leveraging this structured training signal, the MLLM learns to generate format-compliant outputs and grounded reasoning processes as an initial capability, which is critical for subsequent CoT-guided generalization under the GRPO.
@@ 3.3 Group Relative Policy Optimization
Following the Deepseek-R1 [7], the Group Relative Policy Optimization (GRPO) is adopted as the reinforcement learning (RL) algorithm for optimizing our Relation-R1 framework, as illustrated in Figure 2(b). Distinct from critic-dependent methods [32], GRPO directly calculates advantage estimates through response group comparisons sampled from the policy model, inherently eliminating the need for an extra critic network and reducing computational complexity [7]. Key components include the a question $q$, policy model $\pi_{\theta}$, response group $\{o_{i}\}$, and a frozen reference model $\pi_{\theta_{\text{ref}}}$, which acts as a regularization baseline to preserve prior knowledge. The GRPO objective function $J_{\text{GRPO}}(\theta)$ balances reward maximization and policy stability by leveraging following objectives:
$J_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{Q},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}$ (1)
$\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\rho_{i}A_{i},\text{clip}(\rho_{i},1-\epsilon,1+\epsilon)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\right],$
@@
$A_{i}=\frac{r_{i}-\text{mean}\left(\{r_{1},\ldots,r_{G}\}\right)}{\text{std}\left(\{r_{1},\ldots,r_{G}\}\right)},$ (2)
here, $r_{i}$ denotes the reward for response $o_{i}$. The KL divergence term enforces proximity to the reference policy is formulated as:
$D_{KL}(\pi_{\theta}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\left(\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}\right)-1,$ (3)
which ensures controlled exploration without excessive divergence from $\pi_{\theta_{\text{ref}}}$.
The reward function $r_{i}$ is formulated to align with the model’s learning objective for the unified relation understanding. It incorporates rule-based rewards, which consist of three crucial components: Format Compliance Rewards, Binary Relation Rewards, and $N$-nary Relation Rewards. The former enforces syntactic validity by penalizing outputs that deviate from predefined structural constraints. The latter two components focus on task-specific performance.
@@
$r_{\text{form}}(o_{i})=\begin{cases}1&\text{if }o_{i}\text{ adheres to the format},\\ 0&\text{otherwise}.\end{cases}$ (4)
@@
Binary Relation Rewards. Inspired by established evaluation practices [4, 1], we devise the sample-level triplet Recall and sample-level mean Recall (mRecall) for each predicate category in the response of each sample. Specifically, a triplet is deemed correct if:
  • Triplet Accuracy: All entity categories (subject, predicate, object) in this sample are correctly identified.
@@
$r_{\text{binary}}(o)=\alpha\cdot Recall+(1-\alpha)\cdot mRecall,$ (5)
@@
N-ary Relation Rewards. Similarly, to evaluate higher-order relational structures in grounded situation frames, we introduce N-ary Relation Rewards, which assess the model’s ability to detect primary activity and multi-entity interactions. Specifically, we calculate the verb accuracy $Acc_{v}$ and noun accuracy $Acc_{n}$:
  • Verb Accuracy: The predicted primary activity (verb) in the grounded situation frame is the same as the ground-truth activity.
  • Noun Accuracy: All entities involved in the event must satisfy IoU >=0.5 with their ground-truth bounding boxes, and their roles must align with the event’s semantic structure.
@@
$r_{\text{n-ary}}(o)=\beta\cdot Acc_{v}+(1-\beta)\cdot Acc_{n},$ (6)
where hyperparameter $\beta$ balances the trade-off between event type precision and multi-entity structural accuracy. During training, we design a multi-task gate that dynamically selects task-specific rewards based on the presence of <ref> tags in the solutions.

### v2

# Lin Li1,2∗,  Wei Chen1∗,  Jiahui Li3,  Kwang-Ting Cheng1,2  Long Chen1†

@@ Abstract
Recent advances in multi-modal large language models (MLLMs) have significantly improved object-level grounding and region captioning. However, they remain limited in visual relation understanding, struggling even with binary relation detection, let alone N-ary relations involving multiple semantic roles. The core reason is the lack of modeling for structural semantic dependencies among multi-entities, leading to unreliable outputs, hallucinations, and over-reliance on language priors (e.g., defaulting to “person drinks a milk” if a person is merely holding it). To this end, we propose Relation-R1, the first unified relation comprehension framework that explicitly integrates cognitive chain-of-thought (CoT)-guided supervised fine-tuning (SFT) and group relative policy optimization (GRPO) within a reinforcement learning (RL) paradigm. Specifically, we first establish foundational reasoning capabilities via SFT, enforcing structured outputs with thinking processes. Then, GRPO is utilized to refine these outputs via multi-rewards optimization, prioritizing visual-semantic grounding over language-induced biases, thereby improving generalization capability. Furthermore, we investigate the impact of various CoT strategies within this framework, demonstrating that a specific-to-general progressive approach in CoT guidance further improves generalization, especially in capturing synonymous N-ary relations. Extensive experiments on widely-used PSG and SWiG datasets demonstrate that Relation-R1 achieves state-of-the-art performance in both binary and N-ary relation understanding.
@@ 1 Introduction
Recent advancements in MLLMs have significantly enhanced holistic image understanding [1, 2] and object-level grounding [3, 4, 5, 6, 7] capabilities, enabling tasks like region captioning and referring question answering. However, current MLLMs exhibit critical limitations in relational scene understanding — a core competency for achieving human-like visual cognition, as illustrated in Figure 1(a).
@@
Despite notable progress in binary relation detection, i.e., identifying subject-object interactions (e.g., “child-drinking-glass”), current models [8, 9] struggle to capture N-nary relationships. Accurately modeling such relationships requires models to identify multiple N-th entities engaged in the relation with their distinct semantic roles (what, who, and where concepts) [10, 11]. The neglect of these crucial structural semantic dependencies among multi-entities, such as the glass’s role as the functional container of milk in the drinking action, may result in suboptimal relation triplets, e.g., “child-drinking-glass”. Furthermore, these unreliable and superficial pair-wise relations may make the MLLMs prone to both hallucinations and over-reliance on language priors rather than visual-semantic cues (e.g., evidence from semantically grounded visual regions). Consequently, there arises an urgent demand for a unified framework capable of addressing these shortcomings. This raises a critical question: How can we devise a unified framework for joint binary and N-ary relational reasoning with grounded cues?
@@
Intuitively, such complicated relation comprehension requires both strong reasoning and generalization abilities. As for reasoning, this capability is essential as complex relations are often implicit and require the model to perform sophisticated inference by integrating diverse and often subtle visual cues. This involves a multi-step reasoning process, e.g., entity identification (e.g., recognizing “glass” category appears in the image), relation recognition (e.g., recognizing “drink” action appears in the image), role inference (e.g., “glass” as a container), localization, and contextual integration (e.g., observing a child holding a glass near their mouth suggests the “drink” action is likely occurring). In terms of generalization, models trained on limited data struggle to extrapolate to novel interactions and entity combinations or ambiguous contexts (e.g., “cup” as a container for “drinking” action). In addition, inductive bias from language priors (i.e., common-sense assumptions learned purely from text that may not always align with the visual reality) often dominates over visual grounding, further limiting their robustness. For example, a model might struggle to recognize “drinking” from “milk” if its language training strongly associates this action with “glass”.
Fortunately, DeepSeek-R1-Zero [12] has successfully demonstrated the emergence of reasoning capabilities in LLMs purely through RL. However, due to the inherent difficulty of reference models in understanding structural relationships, the direct application of RL often struggles to produce consistently formatted outputs that are claimed in the input prompt [13]. Conversely, when solely adopting supervised fine-tuning, the model suffers from poor generalization due to its overfitting to fixed training patterns and the limited diversity of annotated relational data.
To this end, inspired by the breakthrough of R1-series [12, 14, 15] in enabling reasoning and the ability of SFT to provide structured guidance, we propose a unified two-stage relation comprehension framework, Relation-R1, which combines the advantages of both supervised fine-tuning and reinforcement learning to empower MLLMs with relational reasoning and generalization capability:
Stage 1. SFT: Establishes foundational reasoning through step-by-step cognitive chain-of-thought guidance. This CoT enables the model to break down complex relation comprehension into a sequence of smaller, interpretable steps with visual evidence, such as object detection, spatial localization, etc. Thereby, the model learns to think based on grounded visual content over language inductive bias. Furthermore, its annotated relation-contained instructions ensure that the model generates outputs in a standardized format, facilitating the subsequent computation of rewards during RL.
Stage 2. RL: Refines the structured outputs generated by the SFT model through group relative policy optimization. The learning process is guided by multi-rewards: 1) format rewards, which incentivize the generation of thinking process [12]; 2) binary relation rewards, which encourage accurate identification of pairwise relationships between entities; and 3) N-ary relation rewards, which promote the comprehensive understanding of more complex, multi-entity activity. This stage employs policy gradient updates to cultivate the model to explore various potential solutions and optimizes its output based on defined multi-rewards, fostering robust relational reasoning and generalization.
Furthermore, to gain deeper insights into varying cognitive CoTs, we investigate the impact of template-based CoT and MLLM-generated CoT strategies on binary and N-ary relation comprehension, separately. Notably, by employing a progressive paradigm that initially adopts the template-based CoTs to guide the normative cognitive reasoning process, and then fine-tunes the model with the variable reasoning pathways introduced by a few MLLM-generated CoTs, our Relation-R1 demonstrates better generalization ability. Especially in the N-ary relations, the capacity to explore synonymous relational expressions has emerged.
@@
We validate the effectiveness of Relation-R1 on both the binary relation comprehension task (e.g., scene graph generation [16, 8]) and N-ary relation comprehension task (e.g., grounded situation recognition [10]) across diverse experimental settings. Experiments on the PSG [17] dataset and SWiG [10] dataset demonstrate state-of-the-art performance in relational reasoning while maintaining generalization in ambiguous contexts. In summary, our contributions are fourfold:
@@
  • We reveal the limitations of the existing MLLMs in relation understanding, and incorporate both binary and N-ary relation detection into one unified relation comprehension framework.
@@
  • We propose Relation-R1, a unified framework that integrates cognitive CoT-guided SFT with RL. To the best of our knowledge, this is the first work that enables both relational reasoning and robust generalization in a deep-thinking fashion.
  • We comparatively analyze template-based and MLLM-generated CoTs, demonstrating that progressively guiding the learning process with CoTs in a specific-to-general manner enhances generalization capabilities, particularly in capturing synonymous N-ary relations.
@@
  • Experimental results on the prevalent relation understanding datasets (e.g., +6.84$\sim$6.90gains on PSG dataset) under different settings demonstrate the effectiveness of Relation-R1.
@@ 2 Related Work
Reasoning in MLLMs. Recent LLMs have demonstrated advanced reasoning capabilities by emulating human-like stepwise thinking, significantly enhancing performance on complex tasks [18, 12]. A pivotal breakthrough, DeepSeek-R1 [12], leverages large-scale RL with formatting and result-oriented rewards to enable LLMs to autonomously generate human-like complex Chain-of-Thought (CoT) reasoning, achieving state-of-the-art results across diverse domains. MLLMs extend these capabilities through systematic integration of cross-modal knowledge representation and task-specific rewards [14, 19, 20] For multimodal mathematical reasoning, frameworks such as R1-Onevision [14], Virgo [19], and MM-Eureka [21] jointly reason over visual and textual inputs to solve quantitative tasks, while Vision-R1 [20] combines cold-start initialization with GRPO to refine complex CoT reasoning. For fine-grained classification and grounding tasks, Visual-RFT [22] and R1-Omni [15] utilize GRPO-based reinforcement algorithms and verifiable rewards to enhance contextual reasoning. As for the pixel-level understanding task, SegZero [23] employs reinforcement learning to achieve high-resolution semantic segmentation. Despite these advances, existing MLLMs predominantly focus on object-centric recognition rather than unified visual relation understanding, which requires compositional reasoning about geometric, semantic, and functional interactions between objects and contexts. To bridge this gap, Relation-R1 explicitly integrates relational reasoning into MLLMs via RL, enabling systematic inference for both scene interpretation and event prediction.
@@
Grounded Situation Recognition (GSR). GSR extends scene understanding by jointly modeling actions (verbs) and their associated semantic roles. Pratt et al. [10] pioneered this field with a two-stage RNN-based framework: verb detection followed by noun localization. Subsequent works [11, 35, 36, 37] improved this pipeline by integrating transformers and semantic relation modeling, enabling more coherent scene interpretations. Despite these efforts, previous efforts that depend on annotated training samples may face challenges like long-tailed distribution [24, 25, 26], potentially limiting real-world applicability. While LEX [38] realizes training-free zero-shot GSR, it fails to provide an end-to-end framework and exhibits suboptimal performance and computational efficiency due to its reliance on LLMs’ generated descriptions. This paper proposes Relation-R1, a unified framework that bridges the gap between open-ended capabilities and end-to-end training.
@@ 3 Methodology
Figure 2: Overview of the proposed Relation-R1 framework. (a) SFT with progressive guidance of cognitive CoT, which enforces task-specific reasoning processes and format alignment. (b) RL with GRPO, employs rule-based rewards (e.g., format rewards, binary relation rewards, and N-nary relation rewards) to cultivate robust relation reasoning and generalization ability through policy gradient updates.
@@ 3.1 Preliminaries
Binary Relation Detection. This task requires the model to generate an open-ended scene graph caption given an image, explicitly marking semantic relationships between pairwise objects [8]. The output is a triplet-contained caption annotated with formal tags: objects are denoted by <ref> tags followed by their bounding box coordinates in < box> tags to localize the entity in the image, e.g., <ref>person</ref><box>[[$x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$]]</box>. Predicates are enclosed in <pred> tags, followed by references to their subject and object entities via their corresponding bounding boxes.
N-ary Relation Detection. Similarly, this task aims to output a open-ended grounded situation frame [10] structured as a primary activity (i.e., multi-entity involved relation), followed by a sequence of <role> tags along with corresponding bounding box coordinates. Each role explicitly defines the entity’s participation in the activity, e.g., <agent>person</agent><box>[$x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$]</box>.
@@ 3.2 Supervised Fine-Tuning
As discussed in §1, directly applying GRPO encounters challenges in generating consistently formatted scene graph captions and grounded situation frame outputs. To address this, we employ SFT as a foundational stage to establish task-specific output format adherence (c.f., Figure 2(a)).
However, relying solely on the target answer format during SFT risks compromising the model’s underlying reasoning capabilities by encouraging a superficial alignment with formulated structures [13]. To mitigate this potential issue, we introduce the cognitive chain-of-thought, a methodology that integrates stepwise cognitive processes such as object classification, object grounding, and visual relationship inference. These cognitive CoTs are explicitly encapsulated within <think> tags, during SFT. This ensures that the model retains its capacity for multi-step reasoning while learning to produce structured outputs conforming to the desired format.
To streamline the think process and prevent the MLLM from overfitting to specific CoT patterns during SFT, we adopt a specific-to-general progressive training strategy:
Template-based CoT. For the specific learning phase, we devise a fixed, template-based CoT. As illustrated in Figure 2(a), this template comprises a predefined sequence of steps for both binary and N-ary relation detection. To elaborate, for binary relation detection, the template includes: 1) Object Existence (e.g., the objects present ...); 2) Object Localization (e.g., the objects are located ...); and 3) Relation Existence (e.g., the relations present ...). For N-ary relation detection, it involves: 1) Activity Recognition (e.g., the primary activity is ...); 2) Entities and Roles Recognition (e.g., the entities engaged in the activity are...); and 3) Entity Localization (e.g., the entities are located in ...).
MLLM-generated CoT. For the general reasoning process, we design a cognitive CoT generation prompt that explicitly makes an extra strong MLLM (e.g., Qwen 2.5-VL [39]) to generate step-by-step reasoning processes. To be specific, this cognitive CoT generation prompt consists of a task definition, a ground-truth scene graph of the current image, and a CoT generation instruction. The former one makes the MLLM generate the correct task format. Meanwhile, the latter two anchors reasoning content to visual evidence, ensuring that the MLLM’s cognitive CoT aligns with both semantic validity and visual grounding rather than relying on category priors. The detailed cognitive CoT generation prompt is left in the appendix §B.
Progressive CoT Guidance. With a progressive paradigm that involves initial training on easy-to-learn fully template-based CoTs and subsequent fine-tuning on a small number of MLLM-generated CoTs, the MLLM gains the foundational capability to generate format-compliant outputs and grounded reasoning processes, which is critical for GRPO training.
@@ 3.3 Group Relative Policy Optimization
Following the Deepseek-R1 [12], the group relative policy optimization is adopted as the RL algorithm for optimizing our Relation-R1 framework, as illustrated in Figure 2(b). Distinct from critic-dependent methods [40], GRPO directly calculates advantage estimates through response group comparisons sampled from the policy model, inherently eliminating the need for an extra critic network and reducing computational complexity [12]. Key components include the a question $q$, policy model $\pi_{\theta}$, response group $\{o_{1},o_{2},\ldots,o_{G}\}$, and a frozen reference model $\pi_{\theta_{\text{ref}}}$, which acts as a regularization baseline to preserve prior knowledge. The GRPO objective function $J_{\text{GRPO}}(\theta)$ balances reward maximization and policy stability by leveraging the following objectives:
$J_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim\mathcal{Q},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\rho_{i}A_{i},\text{clip}(\rho_{i},1-\epsilon,1+\epsilon)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\right],$ (1)
@@
$\small A_{i}=\frac{r_{i}-\text{mean}\left(\{r_{1},\ldots,r_{G}\}\right)}{\text{std}\left(\{r_{1},\ldots,r_{G}\}\right)},$ (2)
here $r_{i}$ denotes the reward for response $o_{i}$. The KL divergence term enforces proximity to the reference policy is formulated as:
$\small D_{KL}(\pi_{\theta}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\left(\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}\right)-1.$ (3)
This term ensures controlled exploration without excessive divergence from $\pi_{\theta_{\text{ref}}}$.
The reward $r_{i}$ is formulated to align with the model’s learning objective for the unified relation understanding. It incorporates rule-based rewards, which consist of three crucial components: Format compliance rewards, binary relation rewards, and N-nary relation rewards. The former enforces syntactic validity by penalizing outputs that deviate from predefined structural constraints. The latter two components focus on task-specific performance.
@@
$\small r_{\text{form}}(o_{i})=\begin{cases}1&\text{if }o_{i}\text{ adheres to the format},\\ 0&\text{otherwise}.\end{cases}$ (4)
@@
Binary Relation Rewards. Inspired by established evaluation practices [17, 8], we devise the sample-level triplet recall $R$ and sample-level mean recall $mR$ in the response of each sample. The former is the ratio of correctly predicted triplets to the total ground-truth triplets in a sample; the latter is the average recall across all predicate categories. Specifically, a triplet is deemed correct if:
  • Triplet Accuracy: All entity and relation categories (subject, predicate, object) in this sample are correctly identified.
@@
$\small r_{\text{binary}}(o)=\alpha\cdot R+(1-\alpha)\cdot mR,$ (5)
@@
N-ary Relation Rewards. Similarly, to evaluate higher-order relational structures in grounded situation frames, we introduce N-ary relation rewards, which assess the model’s ability to detect primary activity (verb) and multi-entity interactions. Specifically, we calculate the entity value $V_{e}$ and grounded value $V_{grnd}$ for each sample [10]. These respectively represent the correct proportion in predicted entity categories and roles, and the spatial localization of those entities:
  • Entity Accuracy: An entity within an event is considered correct if its predicted category and semantic role align with the ground truth, and the predicted activity matches the ground-truth.
  • Spatial Localization: The predicted bounding box of each role obtains IoU >=0.5 with the ground-truth’s bounding box.
@@
$\small r_{\text{n-ary}}(o)=\beta\cdot V+(1-\beta)\cdot V_{grnd},$ (6)
where the hyperparameter $\beta$ balances the trade-off between the two values. As the reward’s calculation depends on accurate verb prediction, they jointly evaluate primary activity and multi-entity interactions. During training, we design a multi-task gate that dynamically selects task-specific rewards based on the presence of <ref> tags in the solutions.